Silver sulphide and hydrogen gas are combined to form silver solid and hydrogen

sulphide gas. 10.0 g of H2 and 125.0 g of Ag2S are used in reaction.

What quantity of Ag is produced ? 


1
Expert's answer
2022-04-12T11:18:05-0400

Ag2S + H2 = 2Ag + H2S

n (Ag) = 2 x n (Ag2S) = 2 x n (H2)

n = m / M

M (H2) = 2 g/mol

M (Ag2S) = 247.8 g/mol

n (Ag2S) = 125.0 / 247.8 = 0.5 mol

n (H2) = 10.0 / 2 = 5 mol

Ag2S is the limiting reactant.

n (Ag) = 2 x n (Ag2S) = 2 x 0.5 = 1 mol

M (Ag) = 107.9 g/mol

m (Ag) = n x M = 1 x 107.9 = 107.9 g